OpenStudy (anonymous):

Can someone explain how to solve the equation 2 over 3 k + 6 = −12 for k?

OpenStudy (anonymous):

2/3k + 6 = -12 is that the equation?

OpenStudy (anonymous):

well subtract 6 so you get 2/3 k = -18

OpenStudy (anonymous):

and divide both sides by 2/3... k = -18/2/3

OpenStudy (anonymous):

so you get -3 = k